OpenStudy (anonymous):

Find the error. −5(x + 4) + 21 ≥ −3 + 4(x − 8) −5x − 20 + 21 ≥ −3 + 4x − 32 −5x + 1 ≥ 4x − 35 −9x ≥ −36 x ≥ 1

OpenStudy (anonymous):

When dividing by −9, he did not change the ≥ to ≤. He subtracted 4x from both sides when he should have added 5x. He subtracted 1 from both sides when he should have added 36. He did not make a mistake

OpenStudy (jdoe0001):

\(\bf -5(x + 4) + 21 \ge -3 + 4(x - 8)\\\\ \quad \\ -5x - 20 + 21 \ge -3 + 4x - 32\\\\ \quad \\ -5x + 1 \ge 4x - 35\\\\ \quad \\ \color{red}{-9x \ge -36\\\\ \quad \\ x \ge 1}\)

OpenStudy (jdoe0001):

notice anything there?

OpenStudy (anonymous):

Same thing as the equation above. right?

OpenStudy (jdoe0001):

well... what do you mean?

OpenStudy (anonymous):

It's the same as what I put.

OpenStudy (jdoe0001):

ohh yes

OpenStudy (anonymous):

So I didn't notice anything different..

OpenStudy (jdoe0001):

well, true... I didn't say check for typos, I meant, do you notice anything wrong in the simplification?

OpenStudy (anonymous):

the last step will be x<or equal to 4

OpenStudy (anonymous):

bcoz when u cancel negative sign on both the sides...the sign will also chnge...it will become less than..and it will be 4 as 36/9 is 4 not 1

OpenStudy (anonymous):

so, the error is not switching the signs around ?

OpenStudy (jdoe0001):

well, check again, what do you think is the error?

OpenStudy (jdoe0001):

or....try to simplify it yourself, see what you get

OpenStudy (anonymous):

jst check the last two steps...wat is 36 divided by 9?

OpenStudy (anonymous):

1 or 4?wt d o u think?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yes right..and as i said...if negative sign is present on both the sides..it gets cancelled and the sign chnages.dis is a rule...if u cancel negative on both side sign of equality chnges.u have to remeber dis.